OpenStudy (anonymous):

converge or diverge? An = (1-3/n)^2n

OpenStudy (anonymous):

Most likely Converge.

OpenStudy (anonymous):

Can you prove it? Im stuck at lny = ln(1-3n)/(1/2n)

OpenStudy (zarkon):

\[\left(1-\frac{3}{n}\right)^{2n}\] after taking the natural log you will have \[\frac{\ln\left(1-\frac{3}{n}\right)}{\frac{1}{2n}}\] now use L'Hospitals rule